# Particles before symmetry

### v4

@@ 2.2 Gauge theory and vector bundles: the geometry-first formulation
The geometric perspective I want to develop aims to dispense with the principal fibre bundle altogether. In this Section I set out a formulation of gauge theory that proceeds without gauge potentials, principal bundles, or explicit appeal to gauge symmetries.
The analogy with spacetime helps clarify my aim. Consider $(M,g,\Xi_{i})$, where $(M,g)$ is a smooth Lorentzian manifold and the $\Xi_{i}$ are various tensor fields on $M$, i.e. objects living in spaces constructed from the tangent bundle $TM$. The automorphism group of a typical fibre $T_{x}M$ is $O(3,1)$ (or $SO(3,1)$ if orientation is treated as background structure). This group becomes explicit once we introduce orthonormal frames. Yet much can be said about $g$ and $\Xi_{i}$ in a purely geometric, frame-independent manner, without any reference to $SO(3,1)$. And if instead we were to posit a different group acting on TMTM—say $O(2)$ and not SO(3,1)SO(3,1)—a geometrical rationale would be required to justify that action.
In gauge theory, by contrast, an analogous “frame-free” formulation for the behavior of matter remains largely undeveloped (cf. \@BBOPcitep\@BAP\@BBN(Weatherall, 2016; Gomes, 2024)\@BBCP), and the very idea of a geometric interpretation of the groups and their representations—for example, the adjoint action of $SU(2)$ on $\mathbb{C}^{3}$ endowed with an inner product, as opposed to the fundamental representation of SU(3)SU(3)—is seldom raised. We are after a formulation of gauge theories for which these interpretations are transparent.
I will introduce a realisation of the geometry-first formulation of a certain class of gauge theories, which I will call the vector bundle point of view (VB-POV). [^7] Other kinds of theories could also have a geometry-first formulation, e.g. those based on Cartan geometry, but here my focus is on particle physics, for which the relevant value spaces are vector bundles. To motivate the VB-POV from interpretational issues with the PFB-POV, let me recall that the main role of $\varpi$ in $(P,M,G)$ is to coordinate covariant derivatives between different associated vector bundles. But what is the physical status of $\varpi$? \@BBOPcitet\@BAP\@BBNJacobs (2023, p. 41)\@BBCP convincingly argues they don’t have one; he concludes:
@@
I agree with Jacobs that this is an issue and in \@BBOPcitep\@BAP\@BBN(Gomes, 2024)\@BBCP I showed that it can be overcome. The introduction of $(P,M,G)$ is unnecessary if particles that interact are all sections of the same vector bundles or of tensor products of the same vector bundles. Tensor products inherit the same covariant derivative, so parallel transport automatically marches in step. In this case we have at a hand a natural ‘common cause’ for the coordination of covariant derivatives, without the introduction of principal bundles. This is a realisation of the geometry-first aim, from a VB-POV.
In more detail, given two vector bundles, $E,E^{\prime}{}$, a covariant derivative on $E$ will induce a covariant derivative on $E^{\prime}{}$ whenever $E^{\prime}{}$ is equal to a general tensor product involving $E$ and its algebraic dual, $E^{*}$. Given $E$ a vector bundle with covariant derivative $\mathrm{D}$}, and $E^{*}$ its dual, we define, for sections $\kappa\in\Gamma(E)$ and $\xi\in\Gamma(E^{*})$:
@@
So, in contrast to Figure 1 we have, for a single fundamental vector bundle, Figure 2.
@@
Matters of gauge invariance are now also seen under a different lens. Any composite object constructed from the basic dynamical variables—the $\nabla_{n}$ and other tensor fields—will be tensorial, that is, covariant under the corresponding automorphism groups; and any scalar formed from such quantities will be invariant under those groups. In this respect, the description closely parallels that of classical general relativity in modern treatments employing abstract-index notation: such treatments scarcely mention “gauge invariance” or “coordinate invariance”; all they require is that covariance be secured at the ground level. [^10] Upon quantisation, as I argue in \@BBOPcitep\@BAP\@BBN(Gomes, 2025b)\@BBCP, superpositions of states may require relating objects across distinct classical possibilities. That, I contend, is where “gauge fixing”—or, more broadly, what I call representational schemes—enters. Gauge fixings become necessary when a fixed reference across classical states is needed. If such references are physical, they can, incidentally, be used to describe content in a gauge-invariant (or gauge-fixed) manner, in the traditional PFB-POV sense.
This vantage point also reframes the earlier question of whether canonical maps can exist between distinct vector bundles. In the PFB-POV, the natural candidate—Equation (2.1)—is well defined only within a single representation. Matters look different here. Once covariance is secured at the ground level, all vector bundles charged under a given interaction are taken to ascend from a single fundamental bundle. In the cases to be explored below, each such fundamental bundle $E_{n}$ has typical fibre $\mathbb{C}^{n}$ and is equipped with an inner product $\langle\,\cdot\,$$\cdot\,\rangle$ and, where appropriate, a complex orientation (or volume form) $\epsilon$. The various associated bundles then appear not as independently defined objects requiring ad hoc identifications, but as systematic constructions from $E_{n}$ itself. Their mutual relations are fully accounted for by the standard functorial machinery of geometry: tensor products, (anti)symmetrisation, dualisation, projections into tensor factors, contractions, interior and inner products, and so forth. For instance, to contract an element of $E^{n}$ with one of $E^{n*}\wedge E^{n*}\otimes E^{n}$, we can use the interior product, which generally is a map:
@@
One might object that a parallel, representation-theoretic argument for relating associated vector bundles could be mounted, mirroring the geometric one I have just given. Perhaps given arbitrary different representations of the same group, for arbitrary vector representation spaces, there are systematic ways to relate these representation spaces that mirror the ones I displayed above. That may well be true, but it is beside the point. Even if such arguments exist—and I have not found or worked one out, and am skeptical that one exists (for representations that are not faithful and surjective onto $\mathsf{Aut}(V)$)—the virtue of the geometric route is that it trades purely on geometrical language, and so it speaks directly to a community trained in geometry rather than in group and representation theory. The mere availability of a geometric formulation that sidesteps representation theory or other more technical algebraic machinery is already a win. My aim, after all, is to broaden the borders of the subject, making it accessible to different habits of thought.
@@
Still, for all its merit, at first pass the VB-POV may seem too narrow to capture the full menagerie of gauge theories employed in physics. Some theories—those built from the exceptional Lie groups, for example—fall outside its reach. And even when a gauge group $G$ is given, it is often a nontrivial matter to “reverse-engineer” a vector space structure for which $\mathsf{Aut}(E_{x})\simeq G$. How, for instance, does one coax $SO(4)$ out of a space whose typical fibre is $\mathbb{C}^{3}$; or $U(1)$ out of a space whose typical fibre is $\mathbb{C}^{n}$ with $n\neq 1$? I will have much more to say about all this in Section 5.111111The Peter–Weyl theorem guarantees that $U(n)$ admits nontrivial representations on $\mathbb{C}^{m}$, but extracting from this a natural structure on $\mathbb{C}^{m}$ that renders the action geometrically meaningful is anything but straightforward.
@@
For all that, the Standard Model of particle physics fits neatly within the VB-POV. In the PFB-POV in which the Standard Model is usually described, every particle field is a section of an associated bundle for a principal fibre bundle whose structure group is $SU(3)\times SU(2)\times U(1)$, and the representations that appear are tensor products of the fundamental representation of these component subgroups ($SU(n)$ or $U(n)$ for appropriate $n$). The VB-POV alternative is available because under any representation of $U(n)$, the corresponding associated bundles can just as well be constructed by geometric means from the fundamental vector bundle—via tensor and exterior products, (anti)symmetrization, determinants, and the like. The VB-POV alternative is compelling (as I will expand on in Section 5), because the particular combination of representation, groups, and vector spaces is particularly suited for a geometrical interpretation. In such cases, a covariant derivative on a single vector bundle suffices to encode one fundamental interaction.
Having surveyed both approaches to gauge theory—the symmetry-first PFB-POV and the geometry-first VB-POV—I now turn to the Higgs mechanism. My aim is to present it from within the VB-POV; the more familiar PFB-POV treatment can be found in any standard textbook (see e.g. \@BBOPcite\@BAP\@BBN(Bleecker, 1981, Ch. 10.3)\@BBCP).
@@ 3 The Higgs mechanism in the geometry-first formulation
In the standard presentation (see, e.g., \@BBOPcite\@BAP\@BBN(Tong, 2025, Ch. 2)\@BBCP\@BBOPcite\@BAP\@BBN(Bleecker, 1981, Ch. 10.3)\@BBCP), the Higgs mechanism is usually described in the language of spontaneous symmetry breaking. One is then compelled to invoke stabilisers, quotients of groups, Killing forms on Lie algebras, Goldstone’s theorem, among other tools. In this section I outline an equivalent treatment, the only difference is that it is phrased entirely in the language of vector bundles, which renders the essential structure transparent without appeal to symmetry. It serves as a proof of principle that, in most applications, rerouting our demonstrations to avoid talk of symmetry does not consign us to intractably long alternatives.
@@ 3.1 The Higgs field as a background
Let $(E^{n},M,\mathbb{C}^{n},\langle\cdot,\cdot\rangle_{n})$ be a Hermitian vector bundle over a Lorentzian manifold $(M,g)$, with fibres $E^{n}_{x}\simeq\mathbb{C}^{n}$ and $\langle\cdot,\cdot\rangle_{n}$ an inner product on $E^{n}$, induced by a Hermitian inner product in $\mathbb{C}^{n}$. In this Section we will consider the space of $\nabla_{n}$, the covariant derivatives on $E^{n}$ compatible with the Hermitan structure. We will omit the subscript $n$ when it is understood from context, as it will be in this Section, so for now we take $\kappa\in\Gamma(E)$ and $\|\kappa(x)\|^{2}:=\langle\kappa(x),\kappa(x)\rangle$ (the generalisation to $\kappa\in\Gamma(E^{i}\otimes\cdots E^{j})$ is straightforward). As usual, the norm of a tensor product factorises linearly, i.e. for $\lambda\in\Gamma(TM)$ and $\xi\in\Gamma(E)$, we have:
@@
Thus, when taking norms of tensors that are not spacetime scalars, $g$ is used for the spacetime (abstract) indices; I will leave it implicit, absorbing it into $\langle\,.\,$$\rangle$.
As further bacground structure we will include $\varphi$: a vector-valued spacetime scalar field $\varphi\in\Gamma(E)$ with an everywhere non-zero norm, $\|\varphi(x)\|>0$. Picking out a constant $v>0$ (later to be the minimum of the potential), we write $\|\varphi(x)\|=(H(x)+v)$, for $H\in C^{\infty}(M)$, and get
@@
First, note that $\nabla e_{0}$ does not contain all the information in $\nabla$: the components of $\nabla$ that do not enter this term are destined to remain ‘massless’. Only the components of $\nabla$ that rotate $e_{0}$ into directions orthogonal to $e_{0}$ appear in the quadratic term, while those preserving $e_{0}$ drop out. That is, from (3.4), using $\mathsf{Re}\,\langle\,.\,$$\rangle$ (see footnote 12), we project $\nabla e_{0}$ along $e_{0}$ and its orthogonal complement to find a geometric description of (3.5) as follows.
Viewing the unit section $e_{0}$ as defining a rank-one subbundle $L:=\mathrm{span}\{e_{0}\}\subset E$, we can interpret its orthogonal complement $L^{\perp}$ as the normal bundle to $L$ within $E$. The covariant derivative $\nabla e_{0}$ then defines a one-form with values in $L^{\perp}$; it probes how $L^{\perp}$ curves within $L$. Namely, what we can call a mixed shape operator for $L^{\perp}$: [^13] In differential geometry, given a hypersurface $S\subset M$ with unit normal $n$, the standard shape operator for $S$ is defined as a tensorial operator $K_{n}:TS\rightarrow TS$, namely $K=\nabla n$. Here our operator will mix indices, $K_{n}:TM\rightarrow L^{\perp}$, but it can still be thought of as an operator describing the shape of $L^{\perp}$ within $L$ as one goes around $M$.
@@
i.e. the norm squared of the second fundamental form, or extrinsic curvature, defined by $e_{0}$. In sum, from the kinetic term for the Higgs, assuming $H$ is small compared to $v$, we obtain from (3.5):
$\langle\nabla\varphi,\nabla\varphi\rangle\approx v^{2}\|K_{e_{0}}\|^{2},$ (3.8)
and so a mass-like term for $K_{e_{0}}$.
The operator $\nabla$ merely provides the affine structure of the bundle, while the Higgs field singles out an internal direction $e_{0}$. The dynamics of that affine structure—once the curvature term is included in the Lagrangian (see Equation (A.10))—couple to $v+H$, with $H$ usually assumed small, precisely for the components that rotate $e_{0}$. The resulting mass terms appear for the shape of the surfaces orthogonal to $e_{0}$.
Up to this point, the resulting equation (3.8) is expressed entirely in abstract-index (tensorial) language, and symmetries do not appear. Of course, we are implicitly ‘breaking the symmetry’ by postulating a background field $\varphi$. But note that since any two non-zero Higgs fields define structurally equivalent directions—without further background structures we cannot discern between two such choices—it is the non-zero norm $v$ that is performing the real ‘breaking’.
This concludes the classical, qualitative, account of the ‘mass acquisition’ mechanism for the affine structure using the Higgs as a fixed background structure.

### v5

@@ 2.2 Gauge theory and vector bundles: the geometry-first formulation
The geometric perspective I develop here dispenses with the principal bundle altogether. This section sets out a formulation of gauge theory that proceeds without gauge potentials, principal bundles, or explicit appeal to gauge symmetries.
The analogy with spacetime clarifies my aim. Consider $(M,g,\Xi_{i})$: a smooth Lorentzian manifold $(M,g)$ with various tensor fields $\Xi_{i}$ on $M$, living in spaces constructed from the tangent bundle $TM$. The automorphism group of a typical fibre $T_{x}M$ is $O(3,1)$ (or $SO(3,1)$ if orientation is background structure). This group becomes explicit once we introduce orthonormal frames. Yet we can say much about $g$ and $\Xi_{i}$ in a purely geometric, frame-independent manner, without ever mentioning $SO(3,1)$. And if we were to posit a different group acting on TMTMO(2)O(2), say, rather than SO(3,1)SO(3,1)—we would need a geometric rationale.
In gauge theory, by contrast, an analogous “frame-free” formulation for the behaviour of matter remains largely undeveloped (cf. \@BBOPcitep\@BAP\@BBN(Weatherall, 2016; Gomes, 2024)\@BBCP), and the very idea of a geometric interpretation of the groups and their representations—for example, the adjoint action of $SU(2)$ on $\mathbb{C}^{3}$ endowed with an inner product, as opposed to the fundamental representation of SU(3)SU(3)—is seldom raised. We are after a formulation of gauge theories for which these interpretations are transparent.
I call this realisation the vector bundle point of view (VB-POV). [^8] Other theories could have geometry-first formulations—those based on Cartan geometry, for instance—but my focus is particle physics, where the relevant spaces are vector bundles. To motivate it, recall that the main role of $\varpi$ in $(P,M,G)$ is to coordinate covariant derivatives between associated bundles. But what is the physical status of $\varpi$? \@BBOPcitet\@BAP\@BBNJacobs (2023, p. 41)\@BBCP argues convincingly that it has none:
@@
I agree with Jacobs that this is a problem, and in \@BBOPcitep\@BAP\@BBN(Gomes, 2024)\@BBCP I showed how to solve it: Principal bundles are unnecessary if interacting particles are all sections of the same vector bundles or of their tensor products. Tensor products inherit the same covariant derivative, so parallel transport automatically marches in step. The tensor structure itself provides the ‘common cause’ for coordination.
In more detail, given two vector bundles, $E,E^{\prime}{}$, a covariant derivative on $E$ will induce a covariant derivative on $E^{\prime}{}$ whenever $E^{\prime}{}$ is equal to a general tensor product involving $E$ and its algebraic dual, $E^{*}$. Given $E$ a vector bundle with covariant derivative $\nabla$, and $E^{*}$ its dual, we define, for sections $\kappa\in\Gamma(E)$ and $\xi\in\Gamma(E^{*})$:
@@
The resulting picture, for a single fundamental vector bundle, is summarised in Figure 2.
@@
Matters of gauge invariance are now also seen under a different lens. Any composite object constructed from the basic dynamical variables—the $\nabla_{n}$ and other tensor fields—will be tensorial, that is, covariant under the corresponding automorphism groups; and any scalar formed from such quantities will be invariant under those groups. In this respect, the description closely parallels that of classical general relativity in modern treatments employing abstract-index notation: such treatments scarcely mention “gauge invariance” or “coordinate invariance”; all they require is that covariance be secured at the ground level. [^11] Upon quantisation, as I argue in \@BBOPcitep\@BAP\@BBN(Gomes, 2025a)\@BBCP, superpositions of states may require relating objects across distinct classical possibilities. That, I contend, is where “gauge fixing”—or, more broadly, what I call representational schemes—enters. Gauge fixings become necessary when a fixed reference across classical states is needed. If such references are physical, they can, incidentally, be used to describe content in a gauge-invariant (or gauge-fixed) manner, in the traditional PFB-POV sense.
This vantage point also reframes the earlier question of whether canonical maps can exist between distinct vector bundles. In the PFB-POV, the natural candidate—Equation (2.1)—is well defined only within a single representation. Matters look different here. Once covariance is secured at the ground level, all vector bundles charged under a given interaction are taken to ascend from a single fundamental bundle. In the cases to be explored below, each such fundamental bundle $E_{n}$ has typical fibre $\mathbb{C}^{n}$ and is equipped with an inner product $\langle\,\cdot\,$$\cdot\,\rangle$ and, where appropriate, a complex orientation (or volume form) $\epsilon$. The various associated bundles then appear not as independently defined objects requiring ad hoc identifications, but as constructions from $E_{n}$ itself. Their mutual relations are fully accounted for by the standard functorial machinery of geometry. For instance, to contract an element of $E^{n}$ with one of $E^{n*}\wedge E^{n*}\otimes E^{n}$, we can use the interior product, which generally is a map:
@@
One might object that a parallel representation-theoretic argument could be mounted. Perhaps there are general ways to relate arbitrary representation spaces that mirror the geometric ones. That may be true, but it is beside the point. The virtue of the geometric route is that it trades purely on geometrical language, speaking directly to those trained in geometry rather than representation theory. The mere availability of a formulation that sidesteps algebraic machinery is already a win.
@@
Still, at first pass the geometry-first picture may seem too narrow to capture the full menagerie of gauge theories. Some theories—those built from exceptional Lie groups—seem to fall outside its reach. And even when a gauge group $G$ is given, it can be nontrivial to reverse-engineer a vector space for which $\mathsf{Aut}(E_{x})\simeq G$. [^12] The Peter–Weyl theorem guarantees that $U(n)$ admits nontrivial representations on $\mathbb{C}^{m}$, but extracting a natural structure on $\mathbb{C}^{m}$ that renders the action geometrically meaningful may not be straightforward.
@@
For all that, the Standard Model fits neatly within the geometry-first picture. In the standard PFB-POV every particle field is a section of an associated bundle for a principal bundle with structure group $SU(3)\times SU(2)\times U(1)$, and the representations that appear are tensor products of fundamental representations. The geometry-first alternative is available because the corresponding associated bundles can be constructed geometrically from fundamental vector bundles—via tensor and exterior products, (anti)symmetrisation, determinants, and the like. Namely, we fix once and for all the three fundamental vector bundles (cf. \@BBOPcite\@BAP\@BBNGomes (2024)\@BBCP):
$\big(E^{3},M,\mathbb{C}^{3},\langle\cdot,\cdot\rangle_{3},\epsilon_{3}\big),\qquad\big(E^{2},\mathbb{C}^{2},\langle\cdot,\cdot\rangle_{2},\epsilon_{2}\big),\qquad\big(E^{1},\mathbb{C},\langle\cdot,\cdot\rangle_{1}\big).$ (2.10)
I will refer to (2.10) as the fundamental bundle data. The gauge group in the VB-POV is then
$G_{\mathrm{VB}}:=\mathsf{Aut}(E^{3},\langle\cdot,\cdot\rangle_{3},\epsilon_{3})\times\mathsf{Aut}(E^{2},\langle\cdot,\cdot\rangle_{2},\epsilon_{2})\times\mathsf{Aut}(E^{1},\langle\cdot,\cdot\rangle_{1})\cong SU(3)\times SU(2)\times U(1),$ (2.11)
where each factor is the structure-preserving automorphism group of the corresponding fibre. A covariant derivative on a single vector bundle suffices to encode one fundamental interaction.
With both formulations in hand, the natural next step is to put the VB-POV to work. I begin with the Higgs mechanism; the more familiar PFB-POV treatment can be found in any standard textbook (see e.g. \@BBOPcite\@BAP\@BBN(Bleecker, 1981, Ch. 10.3)\@BBCP).
@@ 3 The Higgs mechanism in the geometry-first formulation
In the standard presentation (see, e.g., \@BBOPcite\@BAP\@BBN(Tong, 2025, Ch. 2)\@BBCP\@BBOPcite\@BAP\@BBN(Bleecker, 1981, Ch. 10.3)\@BBCP), the Higgs mechanism is usually described in the language of spontaneous symmetry breaking. One is then compelled to invoke stabilisers, quotients of groups, Killing forms on Lie algebras, Goldstone’s theorem, among other tools. In this section I outline an equivalent treatment, differing only in that it is phrased entirely in the language of vector bundles; the essential structure becomes visible without appeal to symmetry. It serves as a proof of principle that, in most applications, rerouting our demonstrations to avoid talk of symmetry does not consign us to intractably long alternatives (a common objection to some well-known reformulations; e.g. the Einstein algebra formulation of general relativity \@BBOPcitep\@BAP\@BBN(Geroch, 1972)\@BBCP).
@@ 3.1 The Higgs field as a background
Let $(E^{n},M,\mathbb{C}^{n},\langle\cdot,\cdot\rangle_{n})$ be a Hermitian vector bundle over a Lorentzian manifold $(M,g)$, with fibres $E^{n}_{x}\simeq\mathbb{C}^{n}$ and $\langle\cdot,\cdot\rangle_{n}$ an inner product on $E^{n}$, induced by a Hermitian inner product in $\mathbb{C}^{n}$. In this Section we will consider the space of $\nabla_{n}$, the covariant derivatives on $E^{n}$ compatible with the Hermitian structure. We will omit the subscript $n$ when it is understood from context, as it will be in this Section, so for now we take $\kappa\in\Gamma(E)$ and $\|\kappa(x)\|^{2}:=\langle\kappa(x),\kappa(x)\rangle$ (the generalisation to $\kappa\in\Gamma(E^{i}\otimes\cdots E^{j})$ is straightforward). As usual, the norm of a tensor product factorises linearly, i.e. for $\lambda\in\Gamma(TM)$ and $\xi\in\Gamma(E)$, we have:
@@
Thus, when taking norms of tensors that are not spacetime scalars, $g$ is used for the spacetime (abstract) indices, but I will leave it implicit, absorbing it into $\langle\,.\,$$\rangle$.
As further background structure we will include $\varphi$: a vector-valued spacetime scalar field $\varphi\in\Gamma(E)$ with an everywhere non-zero norm, $\|\varphi(x)\|>0$. [^13] This construction extends smoothly when $\varphi$ is only almost everywhere non-vanishing: define $e_{0}=\varphi/\|\varphi\|$ on the open set where $\varphi\neq 0$, and extend the relevant fields smoothly across the zero set under the usual regularity assumptions. Picking out a constant $v>0$ (later to be the minimum of the potential), we write $\|\varphi(x)\|=(H(x)+v)$, for $H\in C^{\infty}(M)$, and get
@@
First, note that $\nabla e_{0}$ does not contain all the information in $\nabla$: the components of $\nabla$ that do not enter $\nabla e_{0}$ are destined to remain ‘massless’. From (3.4), only some of the components of $\nabla$ that rotate $e_{0}$ into directions orthogonal to $e_{0}$ appear in the quadratic term, while those preserving $e_{0}$ drop out. In more detail, from (3.4), using $\mathsf{Re}\,\langle\,.\,$$\rangle$ (see footnote 14), we project $\nabla e_{0}$ along $e_{0}$ and its orthogonal complement to find a geometric description of (3.5) as follows.
Viewing the unit section $e_{0}$ as defining a rank-one subbundle $L:=\mathrm{span}\{e_{0}\}\subset E$, we can interpret its orthogonal complement with respect to $\text{Re}\langle\cdot,\cdot\rangle$ $L^{\perp}$ as a normal bundle to $L$ within $E$. The covariant derivative $\nabla e_{0}$ then defines a one-form with values in $L^{\perp}$; it probes how $L^{\perp}$ curves within $L$. Namely, what we can call a mixed shape operator for $L^{\perp}$: [^15] In differential geometry, given a hypersurface $S\subset M$ with unit normal $n$, the standard shape operator for $S$ is defined as a tensorial operator $K_{n}:TS\rightarrow TS$, namely $K=\nabla n$. Here our operator will mix indices, $K_{n}:TM\rightarrow L^{\perp}$, but it can still be thought of as an operator describing the shape of $L^{\perp}$ within $L$ as one goes around $M$.
@@
i.e. the norm squared of the second fundamental form, or extrinsic curvature, defined by $e_{0}$.
The operator $\nabla$ merely provides the affine structure of the bundle, while the Higgs field singles out an internal direction $e_{0}$. From the kinetic term, assuming $H$ small compared to $v$, we obtain:
$\langle\nabla\varphi,\nabla\varphi\rangle\approx v^{2}\|K_{e_{0}}\|^{2}.$ (3.8)
That is, the affine structure ‘acquires mass’ precisely in the directions orthogonal to the Higgs; the mass is proportional to the vacuum expectation value $v$ and to the extrinsic curvature of the corresponding orthogonal sub-bundle. Note that Equation (3.8) is expressed entirely in abstract-index (tensorial) language, with no appeal to symmetries or their breaking. We are, of course, ‘breaking the symmetry’ by postulating a background field $\varphi$. But since any two non-zero Higgs fields define structurally equivalent directions—without further background structures we cannot discern between two such choices—it is the non-zero norm $v$ that is doing the real work.
This concludes the classical, qualitative, account of ‘mass acquisition’ for the affine structure using the Higgs as a fixed background structure. Two points of contact with the standard formalism should be noted. First, ‘mass acquisition’ is, at this stage, a statement about the quadratic form (3.8), not yet about gauge bosons in the particle-physics sense. The identification of specific massive and massless components requires writing $\nabla={\mathrm{d}}+\omega$ in a frame adapted to $e_{0}$; this is carried out in Appendix B. Second, the decomposition above is local: it presupposes that $\varphi$ is nowhere vanishing. On nontrivial backgrounds the global picture requires patching, but the geometric content of (3.8) is unaffected.